tali78:

Given a line with slope = 7 and passing through (2,−9) , which of the following is the correct point-slope equation of the line? y+9=7(x−2) y−9=7(x+2) y−9=7(x−2) y+9=7(x+2)

mxddi3:

Hmm, okay so remember that point-slope form is: \[y-y1=m(x-x1)\] Remember that m=slope of the line, and we get the x1 and y1 from our ordered pair. x1=2 and y1= -9 Here, our ordered pair is (2,−9) and our slope (m) is 7. so, now plug those numbers into the equation. To help you, we are going to eliminate B and D because it says "m(x-x1)" so we should have an ending of -2 since it says "x-x1" hopefully this helps some, if not, just lmk and i can re-explain